OpenStudy (anonymous):

find the measure of all angles of triangle KLM for K(-2,1), L(3,2), and M(1,-3)

OpenStudy (wolf1728):

Okay, I just made a graphic

OpenStudy (anonymous):

I converted the points into vectors vector KL=(5) LM=(-2) km=(3) (1) (-5) (-4)

OpenStudy (jdoe0001):

am I wrong in assuming this is for trig?

OpenStudy (wolf1728):

I'm not very familiar with those kind of vectors. I'm used to dealing with vectors in physics 45° mag 20 - that kind of thing This is definitely trig

OpenStudy (wolf1728):

Line KL² = 5² + 1² = 26 Line KL = sqrt(26) = 5.099019 I could figure out the othe sides and then the angles but I was wondering if there's an easier way to do this.

OpenStudy (jdoe0001):

heheh, I wasn't thinking on using vectors, I was thinking on using the distance formula to get the sides length and then use the cosines rule

OpenStudy (wolf1728):

Well okay, I was going to find the other 2 sides and then use the trig.

OpenStudy (wolf1728):

Line KM² = 3² + 4² Line KM² = 25 Line KM = 5

OpenStudy (wolf1728):

Line LM² = 5² + 2² Line LM² = 29 Line LM = 5.38516

OpenStudy (wolf1728):

Angle KML = arcsine (3/5) PLUS arcsine(2/5.38516) = 36.87° + 21.801° = 58.671°

OpenStudy (wolf1728):

Angle KLM = 90° -[arcsine(2/5.38516) + arcsine(1/5.0990919)] Angle KLM = 90° - [21.801° + 11.31°] Angle KLM = 56.889°

OpenStudy (wolf1728):

Angle LKM = arcsine (1/5.0990919) PLUS arcsine (4/5) = 11.31° + 53.13° = 64.44° And the angles all sum to 180°